# Bi-Factorial Preference Optimization: Balancing Safety-Helpfulness in Language Models

### v1

@@ Abstract
Fine-tuning large language models (LLMs) on human preferences, typically through reinforcement learning from human feedback (RLHF), has proven successful in enhancing their capabilities. However, ensuring the safety of LLMs during the fine-tuning remains a critical concern, and mitigating the potential conflicts in safety and helpfulness is costly in RLHF. To address this issue, we propose a supervised learning framework called Bi-Factorial Preference Optimization (BFPO), which re-parameterizes a joint RLHF objective of both safety and helpfulness into a single supervised learning objective. In the supervised optimization, a labeling function is used to capture global preferences ranking to balance both safety and helpfulness. To evaluate BFPO, we develop a benchmark including comprehensive discriminative and generative tasks for helpfulness and harmlessness. The results indicate that our method significantly outperforms existing approaches in both safety and helpfulness. Moreover, BFPO eliminates the need for human prompting and annotation in LLM fine-tuning while achieving the same level of safety as methods that heavily rely on human labor, with less than 10% of the computational resources. The training recipes and models will be released.
@@ 1 Introduction
Fine-tuning the large language models (LLMs) on human preferences, also known as model alignment, has been shown to greatly enhance their text generation abilities [1, 2, 3]. When prompted with well-formatted instructions, these models can generate helpful responses for various tasks like scientific questions answering [4], creative writing [5], coding [6, 7], and planning [8, 9]. However, with this instruction following ability, the models can also be prompted to generate harmful content, such as sexist or racist comments, suggestions for committing crimes, or sensitive medical advice [10, 11, 12]. Therefore, it is crucial to ensure that LLMs are both helpful and safe [1, 2, 13].
@@
The challenge of improving both safety and helpfulness of LLMs arises from the inherent tension between these two objectives [13, 14, 15]. A perfectly safe model may refuse to answer even non-harmful questions (Figure 1-left, b), while a highly helpful model (Figure 1-right, a) may compromise safety. Additionally, training a model on a mix of data annotated with different preference objectives can lead to issues, as shown with model (c) in Figure 1, where the model refuses to generate code (left) yet still produces harmful content (right).
To resolve this problem, recent studies propose to train separate reward models tailored to each objective, safety and helpfulness, and optimize LLMs via multi-objective RLHF, which aggregates reward scores over all objectives [13, 14, 16, 17]. However, developing a safety reward model requires a sufficient number of unsafe responses specific to the model being trained, which is both labor-intensive and computationally demanding [14, 17]. In contrast, supervised optimization, built upon re-parameterizing the RL objective into supervised loss, is more efficient in fine-tuning LLMs on human preferences [18, 19, 20]. However, current work typically focuses on re-parameterizing single reward RLHF objective within the supervised learning framework, and extending this re-parameterization to the multi-reward case is not straightforward [21].
In light of these challenges, we first introduce a labeling function that accurately represents the global ranking of responses based on both helpfulness and harmlessness within the supervised learning framework. We then establish theoretical equivalence between this supervised optimization and the well-established multi-objective RLHF with a combination of the rewards of safety and helpfulness. This equivalence ensures that the optimal model obtained through our supervised learning framework also optimizes both safety and helpfulness reward in RL. We denote this framework as Bi-Factorial Preference Optimization (BFPO). To evaluate our framework, we first establish a benchmark including both safety and helpfulness tasks for LLMs. Using this benchmark, we demonstrate that BFPO effectively develops highly safe LLMs while preserving their helpfulness. Our approach relies only on publicly available datasets, and achieves results comparable to those of methods requiring extensive human labeling efforts. Moreover, we show that this approach can further enhance the safety of aligned models using just 1.5K red teaming prompts, entirely eliminating the need for human annotation. Our contributions are summarized as:
@@
  • Using our algorithm, we efficiently improve the harmlessness of open-sourced models by 15% with a public dataset and by 13% with only 1.5K red teaming data, all while preserving helpfulness. Our method achieves safety scores comparable to those of labor-intensive methods without requiring human prompting or annotations.
@@ 2 Preliminary
Notation and Terminology. Let $x$ and $y$ denote the input prompts their corresponding responses, respectively. For any two responses, $y,y^{\prime}$ generated from a prompt $x$, a binary preference label $I(y\succ y^{\prime}|x)$ is provided by a human annotator to whether $y$ is preferred over $y^{\prime}$. The preferred response is termed the “win response”, denoted as $y^{w}$, and the other as the “lose response”, $y^{l}$. A dataset $D=\{(x,y,y^{\prime},I(y\succ y^{\prime}|x))\}$ that contains prompts, multiple responses, and the human preferences over the responses is referred to as a preference dataset.
Following prior art [19], we define the ground-truth preference $p^{*}$ between two responses $y,y^{\prime}$ as the expected preference label across a broad group of human annotators, i.e., $p^{*}(y\succ y^{\prime}|x)=\mathbb{E}\big{[}I(y\succ y^{\prime}|x)\big{]}$. The ground-truth score of a single response $y$ is then the expected value of its paired preferences with all other responses, i.e., $p^{*}(y|x)=\mathbb{E}_{y^{\prime}}\big{[}p^{*}(y\succ y^{\prime}|x)\big{]}$.
RLHF. RLHF typically consists of two phases [22, 23]: supervised reward learning and policy optimization through reinforcement learning (RL). The training of the reward model $r_{\phi}$, parameterized by $\phi$, is framed by Bradley-Terry (BT) modeling [24], which employs the logistic loss to maximize the distance between the output reward scores of win and lose responses,
$\footnotesize\mathcal{L}_{r}(\phi)=-\mathbb{E}_{(x,y^{w},y^{l})\sim D}\big{[}\log\sigma(r_{\phi}(x,y^{w})-r_{\phi}(x,y^{l}))\big{]},$ (1)
where $\sigma$ is a sigmoid function, and $D$ is a preference dataset. The trained reward model $r_{\phi}$ then provides reward scores for the RL phase. The language model $\pi_{\theta}$, or policy in the RL phase, is optimized with the objective of maximizing the KL-regularized reward [25], i.e.,
$\footnotesize\max_{\pi_{\theta}}\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y|x)}\big{[}r_{\phi}(x,y)-\tau\text{KL}\left[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right]\big{]},$ (2)
@@
Multi-objective RLHF. In multi-objective RLHF, Equation 2 is extended to include multiple reward functions, each corresponding to a specific objective [14, 16, 21, 26, 27],
$\footnotesize\max_{\pi_{\theta}}\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y|x)}\big{[}g(r_{\phi_{1}}(x,y),\dots,r_{\phi_{n}}(x,y))-\tau\text{KL}\left[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right]\big{]}\big{]},$ (3)
@@
Direct Preference Optimization (DPO). DPO [18] reveals that the reward $r$ can be re-parameterized in terms of the policy $\pi$, allowing the policy to be optimized through supervised reward learning:
$\footnotesize\min_{\theta}-\mathbb{E}_{(x,y^{w},y^{l})\sim D}\Big{[}\log\sigma\big{(}\tau\log\frac{\pi_{\theta}(y^{w}|x)}{\pi_{\text{ref}}(y^{w}|x)}-\tau\log\frac{\pi_{\theta}(y^{l}|x)}{\pi_{\text{ref}}(y^{l}|x)}\big{)}\Big{]}.$ (4)
@@
Generalization of DPO. Later work [19, 20] further reveals that a single reward $r$ and the optimal solution $\pi^{*}$ of RLHF in Equation 2 are related by the equation $\pi^{*}(y|x)\propto\pi_{\text{ref}}(y|x)\exp\big{(}\tau^{-1}r(y|x)\big{)}.$ When comparing two responses, $y^{w}$ and $y^{l}$, this relationship yields:
$\footnotesize h_{\pi^{*}}(y^{w},y^{l}):=\log\Big{(}\frac{{\pi^{*}}(y^{w}|x)\pi_{\text{ref}}(y^{l}|x)}{{\pi^{*}}(y^{l}|x)\pi_{\text{ref}}(y^{w}|x)}\Big{)}=\tau^{-1}\big{(}r(y^{w}|x)-r(y^{l}|x)\big{)}$ (5)
Given that this equation holds for the optimal policy $\pi^{*}$, we can directly minimize the difference of the two sides of Equation 5 with a supervised loss $\mathcal{L}$
$\footnotesize\min_{\theta}\mathbb{E}_{(x,y^{w},y^{l})\sim D}\Big{[}\mathcal{L}\big{(}h_{\pi_{\theta}}(y^{w},y^{l}),\tau^{-1}g_{I}(y^{w},y^{l}|x)\big{)}\Big{]},$ (6)
where $g_{I}:\mathbb{R}^{2}\to\mathbb{R}$ is a real-valued label function that approximates the value $r(y^{w}|x)-r(y^{l}|x)$. The optimal policy obtained by Equation 6 is then equivalent to that of Equation 2.
@@ 3 BFPO Framework: Bi-Factorial Preference Optimization
In this section, we aim to extend the supervised learning framework in Equation 6 to improve both safety and helpfulness in LLM alignment. Naively, we could combine the helpfulness and safety datasets, treating safer response in safety dataset and more helpful response in the helpfulness dataset as the win response $y^{w}$ in Equation 6. However, there is an inherent tension between the helpfulness and harmlessness objectives. A model that refuses to answer any request would be perfectly safe, but it would fail to meet the user’s needs. Conversely, a highly responsive model that attempts to address all requests, including potentially harmful ones, may compromise safety in favor of helpfulness [28]. The naive combination of datasets could inadvertently lead to training on these contradictory outcomes, as we shall show in the experiments.
On the other hand, previous work [14, 16] has developed successful multi-objective RLHF methods to resolve this tension, with the objective
$\footnotesize\max_{\pi_{\theta}}\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y|x)}\big{[}g(y|x)-\tau\text{KL}\left[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right]\big{]},$ (7)
where $g(x,y)=g(r_{\text{help}}(x,y),r_{\text{safe}}(x,y))$ is a function that combines the helpfulness reward $r_{\text{help}}(x,y)$ and safety reward $r_{\text{safe}}(x,y)$. Therefore, re-parameterizing Equation 7 to a supervised objective leads to an efficient and effective alignment method. The target objective is:
$\footnotesize\min_{\theta}\mathbb{E}_{(x,y^{hw},y^{hl})\sim D}\Big{[}\mathcal{L}\big{(}h_{\pi}(y^{hw},y^{hl}),\tau^{-1}g_{I}(y^{hw},y^{hl}|x)\big{)}\Big{]},$ (8)
where $y^{hw}$ and $y^{hl}$ are the more helpful and less helpful responses. Similarly to Equation 6, $g_{I}$ is the label function that leverages the safety labels $I_{\text{safe}}(y^{hw}),I_{\text{safe}}(y^{hl})$ to approximate the value $g(y^{hw}|x)-g(y^{hl}|x))$, where $g$ is the global reward function in Equation 7.
@@ 3.1 Empirical Labeling function
In previous single-reward optimization methods [18, 19, 20], $g_{I}(y^{w},y^{l}|x)$ in Equation 6 is typically a positive constant. However, in our case, $g_{I}(y^{hw},y^{hl}|x)$, which approximates the global reward disparity between the more helpful response and the less helpful response, i.e., $g(y^{hw}|x)-g(y^{hl}|x)$, should vary depending on the safety of $y^{hw}$ and $y^{hl}$. For example, in Figure 3, response (a) is more helpful than response (b), and the global reward disparity between (a) and (b) should be positive since both are safe. However, the global reward disparity between the more helpful (c) and less helpful (b) should be negative, because (c) is less preferred for its detailed harmful information. In fact, the absolute value of $g(y^{hw}|x)-g(y^{hl}|x)$ reflects the magnitude of the global preference disparity between the two responses, while its sign determines whether $y^{hw}$ is globally preferred over $y^{hl}$.
To assign label values across various $y^{hw},y^{hl}$ pairs, we first globally rank the responses as illustrated in Figure 3. Our guiding principle is a general preference for safe responses, prioritizing helpfulness only if the responses is safe. We desire the helpful and safe responses like (a) in Figure 3, followed by the acceptable non-helpful but safe responses like (b). We remain neutral toward the harmful but unhelpful responses like (c), and we hate the harmful yet exhaustive (helpful) responses like (d).
Given two responses $y^{hw},y^{hl}$, assuming we have their relative helpfulness ranking, there are four classes of pairs based on their safety, illustrated in Figure 3. For 1 and 2, we prefer the safe and more helpful $y^{hw}$ than the other response, so the signs of the labels should be positive. Similarly, the signs of 3 and 4 should be negative. The preference gap for 1 (Desired vs. Neutral) is larger than for 2, thus the magnitude of the labels should be greater in 1. Likewise, the magnitude of labels of 4 should be greater than that of 3. Consequently, the label value of the four class of pairs should be ordered as 1, 2, 3, and 4. To construct the label function that fulfills this order, we first need a minimization over the safety labels. To ensure a positive label for 2, we require a larger scalar weighting the safety of $y^{hw}$ compared to that of $y^{hl}$. We hypothesize the following form for the label function $g_{I}$:
$\footnotesize g_{I}(y^{hw},y^{hl})=B_{3}(B_{1}I_{\text{safe}}(y^{hw}|x)-I_{\text{safe}}(y^{hl}|x)+B_{2}).$ (9)

### v2

@@ Abstract
Fine-tuning large language models (LLMs) on human preferences, typically through reinforcement learning from human feedback (RLHF), has proven successful in enhancing their capabilities. However, ensuring the safety of LLMs during fine-tuning remains a critical concern, and mitigating the potential conflicts in safety and helpfulness is costly in RLHF. To address this issue, we propose a supervised learning framework called Bi-Factorial Preference Optimization (BFPO), which re-parameterizes a joint RLHF objective of both safety and helpfulness into a single supervised learning objective. In supervised optimization, a labeling function is used to capture the global preferences ranking to balance both safety and helpfulness. To evaluate BFPO, we develop a benchmark that includes comprehensive discriminative and generative tasks for helpfulness and harmlessness. The results indicate that our method significantly outperforms existing approaches in both safety and helpfulness. Moreover, BFPO achieves the same level of safety as methods that heavily rely on human labor with less than 10% of the computational resources and human prompting and annotation process. The training recipes can be found here: https://github.com/wx-zhang/bfpo.
@@ 1 Introduction
Fine-tuning the large language models (LLMs) on human preferences, also known as model alignment, has been shown to greatly enhance their text generation abilities (Ouyang et al., 2022; Askell et al., 2021; OpenAI, 2023). When prompted with well-formatted instructions, these models can generate helpful responses for various tasks like scientific questions answering (Singhal et al., 2023), creative writing (Yuan et al., 2022), coding (Chen et al., 2021; Guo et al., 2024), and planning (Wang et al., 2023; Valmeekam et al., 2023). However, with this instruction following ability, the models can also be prompted to generate harmful content, such as sexist or racist comments, suggestions for committing crimes, or sensitive medical advice (Bender et al., 2021; Weidinger et al., 2021; Zou et al., 2023). Therefore, it is crucial to ensure that LLMs are both helpful and safe (Ouyang et al., 2022; Askell et al., 2021; Bai et al., 2022).
@@
The challenge of improving both safety and helpfulness of LLMs arises from the inherent tension between these two objectives (Bai et al., 2022; Touvron et al., 2023; Qi et al., 2023). A perfectly safe model may refuse to answer even non-harmful questions (Figure 1-left, b), while a highly helpful model (Figure 1-right, a) may compromise safety. Additionally, training a model on a mix of data annotated with different preference objectives can lead to issues, as shown with model (c) in Figure 1, where the model refuses to generate code (left) yet still produces harmful content (right).
To resolve this problem, recent studies propose to train separate reward models tailored to each objective, safety and helpfulness, and optimize LLMs via multi-objective RLHF, which aggregates reward scores over all objectives (Bai et al., 2022; Touvron et al., 2023; Dai et al., 2024; Mu et al., 2024). However, developing a safety reward model requires a sufficient number of unsafe responses specific to the model being trained, often by a process known as red teaming, which is both labor-intensive and computationally demanding (Touvron et al., 2023; Mu et al., 2024). In contrast, Rafailov et al. (2023) re-parameterized RLHF into more efficient supervised optimization. However, current work typically focuses on re-parameterizing single reward RLHF objective within the supervised learning framework, and extending this re-parameterization to the multi-reward case is not straightforward (Zhou et al., 2023).
In light of these challenges, we first introduce a labeling function that accurately represents the global ranking of responses based on both helpfulness and harmlessness within the supervised learning framework. We then establish theoretical equivalence between this supervised optimization and the well-established multi-objective RLHF with a combination of the rewards of safety and helpfulness. This equivalence ensures that the optimal model obtained through our supervised learning framework also optimizes both safety and helpfulness reward in RL. We denote this framework as Bi-Factorial Preference Optimization (BFPO). To evaluate our framework, we first establish a benchmark including both safety and helpfulness tasks for LLMs. Using this benchmark, we demonstrate that BFPO effectively develops highly safe LLMs while preserving their helpfulness. Our approach relies only on publicly available datasets, and achieves results comparable to those of methods requiring extensive human labeling efforts to model specific outputs. Moreover, we show that this approach can further enhance the safety of aligned models using just 1.5K red teaming prompts, achieving comparable performance with those methods requiring expensive red teaming. Our contributions are:
@@
  • Using our algorithm, we efficiently improve the harmlessness of open-sourced models by 15% with a public dataset and by 13% with only 1.5K red teaming data, all while preserving helpfulness. Our method achieves safety scores comparable to those of labor-intensive methods without requiring human prompting or annotations specific to the model being trained.
@@ 2 Preliminary
Notation and Terminology. Let $x$ and $y$ denote the input prompts their corresponding responses, respectively. For any two responses, $y,y^{\prime}$ generated from a prompt $x$, we denote $y$ is preferred over $y^{\prime}$ as $y\succ y^{\prime}$. Then human annotators can provide binary preference labels $I(y\succ y^{\prime}|x)$ on whether $y$ is preferred. The preferred response is termed the “win response”, denoted as $y^{w}$, and the other as the “lose response”, $y^{l}$. A dataset $D=\{(x,y,y^{\prime},I(y\succ y^{\prime}|x))\}$ that contains prompts, multiple responses, and the human preferences over the responses is referred to as a preference dataset.
Following Azar et al. (2024), we define the ground-truth preference $p^{*}$ of $y$ over $y^{\prime}$ as the expected preference label across a broad group of human annotators, i.e., $p^{*}(y\succ y^{\prime}|x)=\mathbb{E}\big{[}I(y\succ y^{\prime}|x)\big{]}$. The ground-truth score of a single response $y$ generated by model $\pi$ is then the expected value of its paired preferences with all other responses, i.e., $p^{*}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y\succ\pi}|x)=\mathbb{E}_{y^{\prime}\sim\pi}\big{[}p^{*}(y\succ y^{\prime}|x)\big{]}$.
RLHF. RLHF typically consists of two phases (Stiennon et al., 2020; Zheng et al., 2023): supervised reward learning and policy optimization by reinforcement learning (RL). The training of the reward model $r_{\phi}$, parameterized by $\phi$, is framed by Bradley-Terry (BT) modeling (Bradley & Terry, 1952), which employs the logistic loss to maximize the distance between the output reward scores of win and lose responses,
$\mathcal{L}_{r}(\phi)=-\mathbb{E}_{(x,y^{w},y^{l})\sim D}\big{[}\log\sigma(r_{\phi}(x,y^{w})-r_{\phi}(x,y^{l}))\big{]},$ (1)
where $\sigma$ is a sigmoid function, and $D$ is a preference dataset. The trained reward model $r_{\phi}$ then provides reward scores for the RL phase. The language model $\pi_{\theta}$, or policy in the RL phase, is optimized with the objective of maximizing the KL-regularized reward (Schulman et al., 2017), i.e.,
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y|x)}\big{[}r_{\phi}(x,y)-\tau\text{KL}\left[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right]\big{]},$ (2)
@@
Multi-objective RLHF. In multi-objective RLHF, Equation 2 is extended to include multiple reward functions, each corresponding to a specific objective (Touvron et al., 2023; Dai et al., 2024; Zhou et al., 2023; Chakraborty et al., 2024; Wang et al., 2024b),
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y|x)}\big{[}g(r_{\phi_{1}}(x,y),\dots,r_{\phi_{n}}(x,y))-\tau\text{KL}\left[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right]\big{]}\big{]},$ (3)
@@
Direct Preference Optimization (DPO)Rafailov et al. (2023) reveals that the reward $r$ can be re-parameterized by the policy $\pi$, and the policy can be optimized through supervised reward learning:
$\min_{\theta}-\mathbb{E}_{(x,y^{w},y^{l})\sim D}\Big{[}\log\sigma\big{(}\tau\log\frac{\pi_{\theta}(y^{w}|x)}{\pi_{\text{ref}}(y^{w}|x)}-\tau\log\frac{\pi_{\theta}(y^{l}|x)}{\pi_{\text{ref}}(y^{l}|x)}\big{)}\Big{]}.$ (4)
@@
Generalization of DPO.  Azar et al. (2024); Tang et al. (2024) further reveals that a single reward $r$ and the optimal solution $\pi^{*}$ of RLHF in Equation 2 are related by the equation $\pi^{*}(y|x)\propto\pi_{\text{ref}}(y|x)\exp\big{(}\tau^{-1}r({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x,y})\big{)}.$ When comparing two responses, $y^{w}$ and $y^{l}$, this relationship yields:
$h_{\pi^{*}}(y^{w},y^{l}):=\log\Big{(}\frac{{\pi^{*}}(y^{w}|x)\pi_{\text{ref}}(y^{l}|x)}{{\pi^{*}}(y^{l}|x)\pi_{\text{ref}}(y^{w}|x)}\Big{)}=\tau^{-1}\big{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r(x,y^{w})-r(x,y^{l})}\big{)}.$ (5)
Details of the relationship are elaborated in Theorem 3.1. As Equation 5 holds for the optimal policy $\pi^{*}$, we can directly minimize the difference of the two sides with a supervised loss $\mathcal{L}$
$\min_{\theta}\mathbb{E}_{(x,y^{w},y^{l})\sim D}\Big{[}\mathcal{L}\big{(}h_{\pi_{\theta}}(y^{w},y^{l}),\tau^{-1}g_{I}(y^{w},y^{l}|x)\big{)}\Big{]},$ (6)
where $g_{I}:\mathbb{R}^{2}\to\mathbb{R}$ is a real-valued label function that approximates the value ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r(x,y^{w})-r(x,y^{l})}$. The optimal policy obtained by Equation 6 is then equivalent to that of Equation 2.
@@ 3 BFPO Framework: Bi-Factorial Preference Optimization
In this section, we aim to extend the supervised learning framework in Equation 6 to improve both safety and helpfulness in LLM alignment. Naively, we could combine the helpfulness and safety datasets, treating safer response in safety dataset and more helpful response in the helpfulness dataset as the win response $y^{w}$ in Equation 6. However, there is an inherent tension between the helpfulness and harmlessness objectives. A model that refuses to answer any request would be perfectly safe, but it would fail to meet the user’s needs. Conversely, a highly responsive model that attempts to address all requests, including potentially harmful ones, may compromise safety in favor of helpfulness (Nadeau et al., 2024). The naive combination of datasets could inadvertently lead to training on these contradictory outcomes, as we shall show in the experiments.
On the other hand,  Touvron et al. (2023); Dai et al. (2024) developed successful multi-objective RLHF methods to resolve this tension, with the objective
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y|x)}\big{[}g(y|x)-\tau\text{KL}\left[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right]\big{]},$ (7)
where $g({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y|x})=g(r_{\text{help}}(x,y),r_{\text{safe}}(x,y))$ is a function that combines the helpfulness reward $r_{\text{help}}(x,y)$ and safety reward $r_{\text{safe}}(x,y)$. Therefore, re-parameterizing Equation 7 to a supervised objective leads to an efficient and effective alignment method. The target objective is:
$\min_{\theta}\mathbb{E}_{(x,y^{hw},y^{hl})\sim D}\Big{[}\mathcal{L}\big{(}h_{\pi}(y^{hw},y^{hl}),\tau^{-1}g_{I}(y^{hw},y^{hl}|x)\big{)}\Big{]},$ (8)
where $y^{hw}$ and $y^{hl}$ are the more helpful and less helpful responses, and as we defined in Equation 5
$h_{\pi}(y^{hw},y^{hl})=\log(\frac{{\pi}(y^{hw}|x)\pi_{\text{ref}}(y^{hl}|x)}{{\pi}(y^{hl}|x)\pi_{\text{ref}}(y^{hw}|x)}),$
and $g_{I}$ is the label function that leverages the safety labels $I_{\text{safe}}(y^{hw}|x),I_{\text{safe}}(y^{hl}|x)$ to approximate the value $g(y^{hw}|x)-g(y^{hl}|x)$, where $g$ is the global reward function in Equation 7.
@@ 3.1 Empirical Labeling function
In previous single-reward optimization methods (Rafailov et al., 2023; Azar et al., 2024; Tang et al., 2024), $g_{I}(y^{w},y^{l}|x)$ in Equation 6 is typically a positive constant. However, in our case, $g_{I}(y^{hw},y^{hl}|x)$, which approximates the global reward disparity between the more helpful response and the less helpful response, i.e., $g(y^{hw}|x)-g(y^{hl}|x)$, should vary depending on the safety of $y^{hw}$ and $y^{hl}$. For example, in Figure 3, response (a) is more helpful than response (b), and the global reward disparity between (a) and (b) should be positive since both are safe. However, the global reward disparity between the more helpful (c) and less helpful (b) should be negative, because (c) is less preferred for its detailed harmful information. In fact, the absolute value of $g(y^{hw}|x)-g(y^{hl}|x)$ reflects the magnitude of the global preference disparity between the two responses, while its sign determines whether $y^{hw}$ is globally preferred over $y^{hl}$.
To assign label values across various $y^{hw},y^{hl}$ pairs, we first globally rank the responses as illustrated in Figure 3. Our guiding principle is a general preference for safe responses, prioritizing helpfulness only if the responses is safe. We desire the helpful and safe responses like (a) in Figure 3, followed by the acceptable non-helpful but safe responses like (b). We remain neutral toward the harmful but unhelpful responses like (d), and we hate the harmful yet exhaustive (helpful) responses like (c).
Given two responses $y^{hw},y^{hl}$, assuming we have their relative helpfulness ranking, there are four classes of pairs based on their safety, illustrated in Figure 3. For 1 and 2, we prefer the safe and more helpful $y^{hw}$ than the other response, so the signs of the labels should be positive. Similarly, the signs of 3 and 4 should be negative. The preference gap for 1 (Desired vs. Neutral) is larger than for 2, thus the magnitude of the labels should be greater in 1. Likewise, the magnitude of labels of 4 should be greater than that of 3. Consequently, the label value of the four class of pairs should be ordered as 1, 2, 3, and 4. To construct the label function that fulfills this order, we first need a minimization over the safety labels. To ensure a positive label for 2, we require a larger scalar weighting the safety of $y^{hw}$ compared to that of $y^{hl}$. We hypothesize the label function $g_{I}$ as:
$g_{I}(y^{hw},y^{hl}|x)=B_{3}(B_{1}I_{\text{safe}}(y^{hw}|x)-I_{\text{safe}}(y^{hl}|x)+B_{2}).$ (9)